at 8 am two cars 315 km apart head toward each other. car a is 20 km/hour faster. at 3pm they meet up. how fast are they?
Answer by mananth(16946)

let speed of car B be x km/h
speed of car A = x+20 km/h
distance = 315 km
time = 7 hours . 8am to 3 pm
since they are travelling towards each other, add up the speeds.
combined speed = x+x+20= 2x+20
d= s * t
315 = 7(2x+20)
315 = 14x+140
14x=315-140
14x= 175
/14
x= 175/14
x= 12.5 km/h speed of car B
speed of car A = 12.5 + 20 = 32.5 km/h